OpenStudy (anonymous):

F(x)=x^3-(a+b)x^2+abc find the value of f(a)

OpenStudy (anonymous):

Plug a in for every x in the function: F(a) = a^3 - (a + b)a^2 + abc F(a) = a^3 - a^3 - ba^2 + abc F(a) = -ba^2 + abc